OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 –6 All real numbers No solution

OpenStudy (noseboy908):

first distribute the -3's

OpenStudy (anonymous):

-3+3 = -3+1-5

OpenStudy (noseboy908):

don't forget the x's, and when you distribute the -3, you have to multiply the 3 (for example) by the-3

OpenStudy (anonymous):

with -6 and 6 they do not equal each other

OpenStudy (anonymous):

-3x+3 = -3x +1 -5

OpenStudy (anonymous):

???

OpenStudy (noseboy908):

one moment, let me draw a picture for you

OpenStudy (anonymous):

okay

OpenStudy (noseboy908):